Make y the subject of the formula:
P = y/y+1
steps, please :)

Answer by Alan3354(69443)

Make y the subject of the formula:
P = y/y+1
P*(y+1) = y
Py + P = y
Py - y = -P
y*(P - 1) = -P
y = -P/(P - 1) = P/(1-P)